40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed May 9 00:45:13 2012, Jenne, Update, CDS, No signals for DTT from SUS OpLevs_OSEMs_allSUSchannels_ADCnoiseOnly.pdfScreenshot.png
    Reply  Wed May 9 01:14:44 2012, Jenne, Update, CDS, No signals for DTT from SUS 
       Reply  Wed May 9 08:21:42 2012, Jamie, Update, CDS, No signals for DTT from SUS 
Message ID: 6627     Entry time: Wed May 9 00:45:13 2012     Reply to this: 6628
Author: Jenne 
Type: Update 
Category: CDS 
Subject: No signals for DTT from SUS 

Upgrades suck.  Or at least making everything work again after the upgrade.

On the to-do list tonight:  look at OSEM sensor and OpLev spectra for PRM, when PRMI is locked and unlocked.  Goal is to see if the PRM is really moving wildly ("crazy" as Kiwamu always described it) when it's nicely aligned and PRMI is locked, or if it's an artifact of lever arm between PRM and the cameras (REFL and AS).

However, I can't get signals on DTT.  So far I've checked a bunch of signals for SUS-PRM, and they all either (a) are just digital 0 or (b) are ADC noise.  Lame.

Steve's elog 5630 shows what reasonable OpLev spectra should look like:  exactly what you'd expect.

Attached below is a small sampling of different SUS-PRM signals.  I'm going to check some other optics, other models on c1sus, etc, to see if I can narrow down where the problem is.  LSC signals are fine (I checked AS55Q, for example).

UPDATE:  SRM channels are same ADC noise.  MC1 channels are totally fine.  And Den had been looking at channels on the RFM model earlier today, which were fine.

ETMY channels - C1:SUS-ETMY_LLCOIL_IN1 and C1:SUS-ETMY_SUSPOS_IN1 both returned "unable to obtain measurement data".  OSEM sensor channels and OpLev _PERROR channel were digital zeros.

ETMX channels were fine

UPDATE UPDATE:  Genius me just checked the FE status screen again.  It was fine ~an hour ago when I sat down to start interferometer-izing for the night, but now the SUS model and both of the ETMY computer models are having problems connecting to the fb.  *sigh* 

Attachment 1: OpLevs_OSEMs_allSUSchannels_ADCnoiseOnly.pdf  48 kB  | Hide | Hide all
OpLevs_OSEMs_allSUSchannels_ADCnoiseOnly.pdf
Attachment 2: Screenshot.png  23 kB  Uploaded Wed May 9 01:58:39 2012  | Hide | Hide all
Screenshot.png
ELOG V3.1.3-